The vapour pressure of ethanol (C2H3OH) and 1-propanol (C3H7OH) at 35°C are


100 mmHg and 37.6 mmhg, respectively. Assuming ideal behaviour, calculate the


partial vapour pressures of ethanol and 1-propanol over a solution, in which the


mole fraction of ethanol is 0.3.

1
Expert's answer
2021-12-06T21:30:01-0500

For ethanol partial pressure = mole fraction ×Vapour pressure = 0.3 × 100 = 30 mmHg.

Now for propanol partial pressure = 0.7×37.6 = 26.32 mmHg